You have a simple circuit with a total resistance of 5 Ω. If a 3 -A current is maintained in that circuit, how much energy (in Joule) is delivered through the circuit in twenty five seconds?

(1125J IS NOT CORRECT ANSWER)
Answer #1

Energy= Power*Time

Power=I^2*R

Power= 3^2*5

Power= 45 watt

Energy= 45*5

Energy= 225 Joule